OpenStudy (anonymous):

1 2/5-2/3 Estimate. Then find the diffference and write it in simplest form

OpenStudy (anonymous):

I got 11/15 as the answer

OpenStudy (mertsj):

My estimate is 1

OpenStudy (anonymous):

Can you show me how you got 1?

OpenStudy (mertsj):

\[1 \frac{2}{5}-\frac{2}{3}=1 \frac{6}{15}-\frac{10}{15}=\frac{21}{15}-\frac{10}{15}=\frac{11}{15}\]

OpenStudy (mertsj):

I thought that 1 2/5 is a little more than 1 and 2/3 is a little less than 1 so if I subtract, my answer should be close to 1